OpenStudy (anonymous):

x = -7y + 34 and x + 7y = 32

OpenStudy (anonymous):

just plug in x for -7y + 34 and simplify the equation.

OpenStudy (anonymous):

-7y + 34 +x +7y = 32 so the -7y and +7y cancel and you are left with 34+x = 32 32-34 = -2 so x = -2